# KAAE: Numerical Reasoning for Knowledge Graphs via Knowledge-aware Attributes Learning

### v1

@@ C.2. Hyperparameters
To ensure the reproducibility of our model, we list the hyperparameters used in KAAE in Table8. These parameters were selected based on the validation splits across all three datasets: US-Cities, Spotify, and Credit.
Table 8. Hyperparameters to reproduce our results.
Hyperparameter US-Cities Spotify Credit
Entity/Relation Dimension $d_{emb}$ 200 200 200
Attribute Dimension $d_{att}$ 200 200 200
Number of Attention Heads L 5 5 5
Number of MoE Experts K 3 3 3
Score Function Bias $\varepsilon$ 9.0 9.0 9.0
Number of Top-K samples 3 3 3
Numerical Reasoning Weight $\Delta$ 0.25 0.25 0.25
CL Loss Coefficient $\lambda$ 0.25 0.25 0.25
Dropout Rate 0.7 0.7 0.7
Learning Rate 0.001 0.001 0.001
Batch Size 200 256 256
Appendix D LEARNING PROCEDURE
@@
The overall learning process of KAAE is shown in Algorithm 1. At the beginning, we initialize the loss. We then use the KA Encoder to obtain the header embedding $\bm{e_{joint,h}^{att}}$ for each triple(h, r, t) in the training set (described in line 5). If r is a numerical comparison relation, we obtain the set $\mathcal{E}^{+}$, $\mathcal{E}^{-}$ of comparative positive and negative samples based on $\bm{e_{joint,h}^{att}}$ (described in lines 7-8). Lines 9-10 show the process of generating high-quality blended positive/negative samples via Top-K sampling. Then, the contrast learning loss is accounted for in the total loss (line 11). Finally, the cross-entropy loss is added according to Eq(14), and finish the updating process.
Algorithm 1 Training KAAE
@@
Appendix E ADDITIONAL EXPERIMENTAL RESULTS
E.1. Feasibility Test
To test the ability of KAAE to learn values based on the magnitude of attribute numerical values, we conducted a feasibility experiment. We convert the values in the Spotify dataset to binary values to distinguish the presence or absence of numeric attributes. We assign a value of 1 to data that exists and a value of 0 to data that does not exist. The result is shown in Table 9. Our proposed model KAAE outperforms the competitors LiteralE(Kristiadi et al., 2019) and RAKGE(Kim et al., 2023). LiteralE did not show significant changes under the two experimental conditions. Both KAAE and RAKGE can compare the magnitude of the values, and KAAE performed better.
Table 9. Feasiblity test on Spotify dataset
Model Setting H@1 H@3 H@10 MR MRR
LiteralE Binary 0.263 0.362 0.480 79 0.339
Numeric 0.266 0.377 0.494 74 0.347
RAKGE Binary 0.394 0.479 0.571 62 0.459
Numeric 0.487 0.602 0.692 37 0.560
KAAE Binary 0.474 0.588 0.683 47 0.550
Numeric 0.636 0.694 0.758 30 0.680
E.2. Sensitivity Experiment
We adjust the hyperparameter $\lambda$ of contrastive learning loss in Eq. (15) to verify the necessity of introducing contrastive learning. At the same time, we changed $\Delta$ in Eq. (13) to test the performance of KAAE with different weights of order-embedding in the score function. The results are shown in Figure 6.
Figure 6. Performance of KAAE with different hyperparameter values of $\lambda$ and $\Delta$ on Credit dataset.
E.3. Loss Experiment
To verify the effectiveness of our proposed two loss functions, we compare the performance of KAAE trained with different loss functions. The results in Table 10 show that both loss functions are reasonable.
(27) $\begin{split}{Loss}_{1}=-\frac{1}{\mathcal{|G|}}\sum\limits_{l\in\mathcal{|G|}}\log\sigma(&score(\bm{e}_{joint,h}^{att},\bm{e}_{r},\bm{e}_{mix}^{t+})\\ -&score(\bm{e}_{joint,h}^{att},\bm{e}_{r},\bm{e}_{mix}^{t-})),\end{split}$
(28) $\begin{split}{Loss}_{2}=-&\frac{1}{\mathcal{|G|}}\sum\nolimits_{l\in\mathcal{G}}\log\frac{exp(\textbf{p}^{\top}\textbf{u}/\tau)}{exp(\textbf{p}^{\top}\textbf{u}/\tau)+\sum\nolimits_{v_{-}\in\bm{e}_{mix}^{t-}}exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)},\end{split}$
Table 10. Loss function test
@@ E.3. Loss Experiment
E.4. Visualization
@@
Figure 7. Visualization of relevance scores of each numeric attribute without/with entity-aware on Credit dataset.
Analysis of Contrastive Learning. To improve the continuity of numerical attributes and the sensitivity of numerical values to very small differences near boundary values, we propose a method to accomplish contrastive learning using ordinal high-quality positive and negative samples. We use the algorithm named T-SNE(Maaten and Hinton, 2008) to visualize the samples in 2D space to observe the distribution of high-quality positive and negative samples. First, we replace the vector $\bm{e_{joint,i}^{att}}$ in Eq.(7), which contains rich information of numerical attribute and order-embedding. Eqs. (10)-(11) are repeated 100 times using a definite attribute value of age-40 with entity and relation to accumulating a large number of tail samples in contrastive learning. We compare KAAE to another well-performed model, RAKGE(Kim et al., 2023). The results in Fig.8 show that our KAAE can improve the generalization ability and robustness of the model by learning difficult samples to get a good representation of entities and make more accurate judgments in numerical reasoning tasks.
@@
(d) KAAE on US-Cities dataset
(e) KAAE on Spotify dataset
(f) KAAE on Credit dataset
Figure 8. Visualization of the entity embeddings using T-SNE and K-means. The first row shows the results of RAKGE on the three datasets and the second row shows the better performance of KAAE. By using KAAE, the nodes are more compact with each other, which means that we can get excellent entity embeddings
E.5. Evaluation on FB15k-237 Dataset
FB15k-237 is a classical dataset used for the standard link prediction task accomplished by knowledge graphs. Therefore, we conducted additional experiments on the FB15k-237 dataset. As we can see from Table11, KAAE is still the state-of-the-art model.
Table 11. Link prediction results on FB15k-237(Toutanova and Chen, 2015). Bold scores represent the best results.
@@ E.5. Evaluation on FB15k-237 Dataset
TransE(Bordes et al., 2013) 0.240 0.368 0.516 182 0.332
DisMult(Dong et al., 2014) 0.202 0.306 0.433 392 0.279
ConvE(Dettmers et al., 2018) 0.232 0.351 0.492 276 0.319
TuckER(Balazevic et al., 2019b) 0.266 0.394 0.544 - 0.358
MuRP(Balazevic et al., 2019a) 0.245 0.370 0.521 - 0.336
R-GCN(Schlichtkrull et al., 2018) 0.153 0.258 0.414 - 0.248
LiterlE(Kristiadi et al., 2019) 0.232 0.348 0.483 280 0.317
KBLRN(García-Durán and Niepert, 2018) 0.215 0.333 0.468 358 0.301
MT-KGNN(Tay et al., 2017) 0.204 0.312 0.445 532 0.285
RAKGE(Kim et al., 2023) 0.282 0.402 0.531 321 0.366
KAAE 0.355 0.464 0.553 164 0.428
E.6. Analysis of Backbones
To validate the compatibility of TransE with KAAE, we replaced different backbones. Specificaly, we changed TransE to RotatE(Sun et al., 2019), SimplE(Kazemi and Poole, 2018) and DistMult. Table12 (in Appendix) shows that our KAAE+TransE model significantly outperforms the KAAE+RotatE model, exhibiting enhancements of 354.2% in Hits@1 and 590.4% in MRR. It also outperforms SimplE and DistMult. We infer that this is because TransE(Bordes et al., 2013) is compatible with sequential embedding, which helps the model to solve the ordering and boundary problems of values and deal with asymmetric relationships. In this sense, replacing the backbones causes the model to have difficulty retaining the ordering information.
E.7. Breakdown Analysis.
@@ E.6. Analysis of Backbones
Based on our current framework, we have selected some of the relations that are difficult to recognize. In dataset US-Cities, the difference in the values of each attribute for each city is almost always less than 1, which makes it difficult to determine the relation is_located_in_county and is_located_in_state. This gives us ideas for improvements. In our future work, we will explore how to target the improvement of the link prediction tasks involving fine-grained numerical size comparison relations. In addition, we are already working the challenges of extreme values on model robustness.
Table 12. Backbone study of KAAE on Credit dataset
Model H@1 H@3 H@10 MR MRR
KAAE+TransE 0.745 0.818 0.888 7 0.794
KAAE+RotatE 0.164 0.116 0.205 151 0.115
KAAE+SimplE 0.026 0.201 0.447 39 0.163
KAAE+DistMult 0.003 0.080 0.200 177 0.070
Relevance of Attributes. Figure 9 illustrates the change in performance of KAAE for different densities of numerical attributes. As the data density decreases, LiteralE, a model that relies on structural information for numerical reasoning, doesn’t change much, at the same time the performance of our KAAE model decreases, but always stays ahead in absolute performance. We can infer that through our MoEKA Encoder and learning by utilizing high-quality positive and negative samples in contrastive learning, KAAE is sensitive to numerical magnitude while possessing the ability to learn adaptively and remain robust despite sparse numerical attributes.
@@
Figure 9. Performance of KAAE with different density of numeric attributes for Spotify dataset. A density of 100% means that all entities have numeric values for every attribute, while 0% means that none of the entities have numeric attributes.

### v3

@@ C.3 Hyperparameter Settings
To ensure the reproducibility of our model, we list the hyperparameters used in NumCoKE in Table 5. These parameters were selected based on the validation splits across all three datasets. All the evaluated models are implemented on a server with one GPU (NVIDIA A800, 80GB).
Hyperparameter Search range US-Cities Spotify Credit
Entity/Relation Dimension $d_{emb}$ [50, 100, 200] 200 200 200
Attribute Dimension $d_{att}$ [50, 100, 200] 200 200 200
Number of Attention Heads L [1, 2, 3, 4, 5] 5 5 5
Number of MoE Experts K [1, 2, 3, 4, 5] 3 3 3
Score Function Bias $\varepsilon$ [1.0, 5.0, 9.0] 9.0 9.0 9.0
Number of Top-K samples [1, 2, 3, 4, 5] 3 3 3
Numerical Reasoning Weight $\Delta$ [0, 0.25, 0.5, 0.75] 0.25 0.25 0.25
CL Loss Coefficient $\lambda$ [0, 0.25, 0.5, 0.75] 0.25 0.25 0.25
Dropout Rate [0.1, 0.3, 0.5, 0.7, 0.9] 0.7 0.7 0.7
Learning Rate [0.0001, 0.0005, 0.001, 0.005] 0.001 0.001 0.001
Batch Size [64, 128, 256, 512] 128 256 256
Table 5: Hyperparameters to reproduce our results.
Appendix D D Learning Procedure (Algorithm)
@@
The overall learning process of NumCoKE is shown in Algorithm 1. At the beginning, we initialize the loss. We then use the KA Encoder to obtain the header embedding $\bm{e_{joint,h}^{att}}$ for each triple(h, r, t) in the training set (described in line 5). If r is a numerical comparison relation, we obtain the set $\mathcal{E}^{+}$, $\mathcal{E}^{-}$ of comparative positive and negative samples based on $\bm{e_{joint,h}^{att}}$ (described in lines 7-8). Lines 9-10 show the process of generating high-quality blended positive/negative samples via Top-K sampling. Then, the contrast learning loss is accounted for in the total loss (line 11). Finally, the cross-entropy loss is added according to Eq(15), and finish the updating process.
Algorithm 1 Training NumCoKE
@@
Appendix E E Supplementary Experiments
E.1 Single Module Ablation Study
We conduct additional ablation experiments on Credit dataset, focusing on only considering the effect of individual modules within NumCoKE (MoE, KA, OS). The experimental hyper-parameters remained the same. As shown in table 6, MoE achieves the greatest improvement (MoE¿OS¿KA), indicating that our designed relation-guided context expert perception plays a crucial role in enabling NumCoKE to focus on and integrate key information.
MoE KA OS H@1 H@3 H@10 MRR
$\times$ $\times$ $\times$ 0.647 0.733 0.823 0.708
$\surd$ $\times$ $\times$ 0.693 0.769 0.849 0.741
$\times$ $\surd$ $\times$ 0.681 0.752 0.837 0.729
$\times$ $\times$ $\surd$ 0.687 0.758 0.846 0.738
$\surd$ $\surd$ $\surd$ 0.745 0.818 0.888 0.794
Table 6: Single module ablation study on Credit dataset
E.2 Hyperparameter Sensitivity Experiment
We adjust the hyperparameter $\lambda$ of contrastive learning loss in Eq. (16) to verify the necessity of introducing contrastive learning. At the same time, we changed $\Delta$ in Eq. (14) to test the performance of NumCoKE with different weights of order-embedding in the score function. The results are shown in Figure 6.
Figure 6: Performance of NumCoKE with different hyperparameter values of $\lambda$ and $\Delta$ on Credit dataset.
E.3 Loss Function Experiment
To verify the effectiveness of our proposed two loss functions for contrastive learning, we compare the performance of NumCoKE trained with different loss functions. The results in Table 7 show that both loss functions are reasonable.
$\begin{split}{Loss}_{1}=-\frac{1}{\mathcal{|G|}}\sum\limits_{l\in\mathcal{|G|}}\log\sigma(&score(\bm{e}_{joint,h}^{att},\bm{e}_{r},\bm{e}_{mix}^{t+})\\ -&score(\bm{e}_{joint,h}^{att},\bm{e}_{r},\bm{e}_{mix}^{t-})),\end{split}$ (28)
$\begin{split}{Loss}_{2}=&-\frac{1}{\mathcal{|G|}}\sum\nolimits_{l\in\mathcal{G}}\log\frac{exp(\textbf{p}^{\top}\textbf{u}/\tau)}{exp(\textbf{p}^{\top}\textbf{u}/\tau)+\sum\nolimits_{v_{-}\in\bm{e}_{mix}^{t-}}exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)},\end{split}$ (29)
@@ E.3 Loss Function Experiment
Table 7: Loss function test.
E.4 Feasibility Test
To test the ability of NumCoKE to learn values based on the magnitude of attribute numerical values, we convert the values in the Spotify dataset to binary values to distinguish the presence or absence of numeric attributes. We assign a value of 1 to data that exists and a value of 0 to data that does not exist. The result is shown in Table 8. Our proposed model NumCoKE outperforms the competitors LiteralE (DBLP:conf/semweb/KristiadiKL0F19) and RAKGE (DBLP:conf/kdd/KimKKPJP23). LiteralE did not show significant changes under the two experimental conditions. Both NumCoKE and RAKGE can compare the magnitude of the values, and NumCoKE performed better.
Model Setting H@1 H@3 H@10 MR MRR
LiteralE Binary 0.263 0.362 0.480 79 0.339
Numeric 0.266 0.377 0.494 74 0.347
RAKGE Binary 0.394 0.479 0.571 62 0.459
Numeric 0.487 0.602 0.692 37 0.560
NumCoKE Binary 0.474 0.588 0.683 47 0.550
Numeric 0.636 0.694 0.758 30 0.680
Table 8: Feasiblity test on Spotify dataset.
E.5 Extreme value Analysis
To test whether our edge-range sampling approach will lead to a weaker perception of large numerical differences, we did an additional experiment to validate NumCoKE by selecting different proportions of positive/negative samples separately on Credit dataset. Table 9 demonstrates that NumCoKE can still maintain a good performance without utilizing all samples.
Proportion H@1 H@3 H@10 MR MRR
20% 0.739 0.805 0.882 7 0.796
40% 0.737 0.802 0.879 7 0.782
60% 0.746 0.821 0.885 7 0.789
80% 0.743 0.812 0.876 7 0.792
100% 0.745 0.818 0.888 7 0.794
Table 9: Performance with different proportions of samples.
Additionally, we added 10% extra extreme samples to the datasets. The results in Table 10 show that extreme values will not affect the performance of NumCoKE.
Dataset Extreme H@1 H@10 MR MRR
US-Cities $\surd$ 0.434 0.643 118 0.505
$\times$ 0.439 0.640 119 0.508
Spotify $\surd$ 0.629 0.763 31 0.678
$\times$ 0.636 0.753 30 0.680
Credit $\surd$ 0.747 0.891 7 0.795
$\times$ 0.745 0.888 7 0.794
Table 10: Performance of NumCoKE with extra extreme values.
E.6 Backbone Model Analysis
To validate the compatibility of TransE with NumCoKE, we replaced different backbones. Specificaly, we changed TransE to RotatE (DBLP:conf/iclr/SunDNT19), SimplE (DBLP:conf/nips/Kazemi018) and DistMult. Table11 shows that our NumCoKE+TransE model significantly outperforms the NumCoKE+RotatE model, exhibiting enhancements of 354.2% in Hits@1 and 590.4% in MRR. It also outperforms SimplE and DistMult. We infer that this is because TransE (DBLP:conf/nips/BordesUGWY13) is compatible with sequential embedding, which helps the model to solve the ordering and boundary problems of values and deal with asymmetric relationships. In this sense, replacing the backbones causes the model to have difficulty retaining the ordering information.
Model H@1 H@3 H@10 MR MRR
NumCoKE+TransE 0.745 0.818 0.888 7 0.794
NumCoKE+RotatE 0.164 0.116 0.205 151 0.115
NumCoKE+SimplE 0.026 0.201 0.447 39 0.163
NumCoKE+DistMult 0.003 0.080 0.200 177 0.070
Table 11: Backbone study of NumCoKE on Credit dataset
E.7 Visualization
@@
Figure 7: Visualization of relevance scores of each numeric attribute without/with entity-aware on Credit dataset.
Analysis of Contrastive Learning. To improve the continuity of numerical attributes and the sensitivity of numerical values to very small differences near boundary values, we propose a method to accomplish contrastive learning using ordinal high-quality positive and negative samples. We use the algorithm named T-SNE (maaten2008visualizing) to visualize the samples in 2D space to observe the distribution of high-quality positive and negative samples. First, we replace the vector $\bm{e_{joint,i}^{att}}$ in Eq.(7), which contains rich information of numerical attribute and order-embedding. Eqs. (10)-(11) are repeated 100 times using a definite attribute value of age-40 with entity and relation to accumulating a large number of tail samples in contrastive learning. We compare NumCoKE to another well-performed model, RAKGE (DBLP:conf/kdd/KimKKPJP23). The results in Fig.8 show that our NumCoKE can improve the generalization ability and robustness of the model by learning difficult samples to get a good representation of entities and make more accurate judgments in numerical reasoning tasks.
@@
(d) NumCoKE on US-Cities dataset
(e) NumCoKE on Spotify dataset
(f) NumCoKE on Credit dataset
Figure 8: Visualization of the entity embeddings using T-SNE and K-means. The first row shows the results of RAKGE on the three datasets and the second row shows the better performance of NumCoKE. By using NumCoKE, the nodes are more compact with each other, which means that we can get excellent entity embeddings
E.8 Evaluation on FB15k-237 Dataset
FB15k-237 is a classical dataset used for the standard link prediction task accomplished by knowledge graphs. Therefore, we conducted additional experiments on the FB15k-237 dataset. As we can see from Table12, NumCoKE is still the state-of-the-art model.
@@ E.8 Evaluation on FB15k-237 Dataset
TransE (2013) 0.240 0.368 0.516 182 0.332
DaBR (2025) 0.202 0.306 0.433 392 0.279
ConvE (2018) 0.232 0.351 0.492 276 0.319
TuckER (2019) 0.266 0.394 0.544 - 0.358
MuRP (2019) 0.245 0.370 0.521 - 0.336
R-GCN (2018) 0.153 0.258 0.414 - 0.248
LiterlE (2019) 0.232 0.348 0.483 280 0.317
KBLRN (2018) 0.215 0.333 0.468 358 0.301
MT-KGNN (2017) 0.204 0.312 0.445 532 0.285
RAKGE (2023) 0.282 0.402 0.531 321 0.366
NumCoKE (Ours) 0.355 0.464 0.553 164 0.428
Table 12: Link prediction results on FB15k-237 (DBLP:conf/acl-cvsc/ToutanovaC15). Bold scores represent the best results.
E.9 Breakdown Analysis
@@
Based on our current framework, we have selected some of the relations that are difficult to recognize. In US-Cities dataset, the difference in the values of each attribute for each city is almost always less than 1, which makes it difficult to determine the relation is_located_in_county and is_located_in_state. This gives us ideas for improvements. In the future, we will explore how to target the improvement of the link prediction tasks involving fine-grained numerical size comparison relations. In addition, we are already working the challenges of extreme values on model robustness.
@@
Figure 9: Performance of NumCoKE with different density of numeric attributes on Spotify. A density of 100% means that all entities have numeric values for every attribute, while 0% means that none of the entities have numeric attributes.
TransE LiteralE RAKGE NumCoKE
Numeric-aware $\times$ $\surd$ $\surd$ $\surd$
Knowledge-aware $\times$ $\times$ $\times$ $\surd$
Relation-aware $\times$ $\times$ $\surd$ $\surd$
Self-Supervised $\times$ $\times$ $\surd$ $\surd$
Mixture of Experts $\times$ $\times$ $\times$ $\surd$
Table 13: Model difference analysis.
Model Head Entity Relation Important Attributes (Attention Score) Predicted Top5 Tail Entities
RAKGE - commute_time_comp income, age, male_population, Citrus Heights, Vineland, Fort Myers,
commute_time(0.0874), limited_english. Farmington Hills, Brooklyn Park.
- ranking_comp hispanic, density, longitude Terre Haute, McGee Creek,
race_white, ranking(0.0732). Lakewood, Decatur, Somerville.
NumCoKE Grand Saline commute_time_comp commute_time(0.291), population, density, Fort Myers, Citrus Heights, Springdale,
family_dual_income, home_value. Brooklyn Park, Deerfield Beach.
Cedarville ranking_comp ranking(0.268), population_proper, widowed, Decatur, Buena Park, Madera
home_valueincome, veteran. Johns Creek, Somerville.
Table 14: Case study on US-Cities dataset. ’-’ means not considering the semantic information from the entity. The text with same background color represents related relations and attributes, the underlined entity indicate the correct answer.
Relevance of Attributes. Figure 9 illustrates the performance change of NumCoKE for different densities of numerical attributes. As the data density decreases, LiteralE, a model that relies on structural information for numerical reasoning, doesn’t change much, at the same time the performance of our NumCoKE model decreases, but always stays ahead in absolute performance. We can infer that through MoEKA Encoder and learning by utilizing high-quality ordinal samples in contrastive learning, NumCoKE is sensitive to numerical magnitude while possessing the ability to learn adaptively and remain robust despite sparse numerical attributes.
Study for Ordinal Knowledge Contrastive Learning. To validate the effectiveness of our ordinal knowledge contrastive learning method, we compared NumCoKE with other self-supervised models. MRR is improved from 0.331 for BiGI and 0.573 for RAKGE to 0.680 on the Spotify dataset. We can deduce that NumCoKE learn richer representations than graph-based approaches, including BiGI, SLiCE, and SimGCL, by considering extra numerical attributes of entities. Compared to RAKGE, NumCoKE focuses on generating and extracting hard, high-quality ordinal samples to determine the ordering among values based on magnitude boundaries and distinguish subtle semantic differences accurately, which makes contrastive learning more effective. Meanwhile, although HAKE, TuckER, and ConvE often outperform TransE, Table 1 suggests otherwise. It is probably because numerical reasoning is a one-to-one task in most cases, and is better suited to TransE. Besides, diffusion model in KGDM may destroy the comparison relation among numerical semantics, resulting in weak performance.
Model difference analysis. We compare NumCoKE with popular models in Table13. Numeric-aware indicates the availability of numeric information, Knowledge-aware indicates Interacting with entities and relations.
E.10 Case Study
We explore how NumCoKE performs better on predicting triples by considering numerical attributes more relevant to the context by investigating the significant attributes that contributed to inference of a derived relation. We design a two-stage demonstration of “Attributes Attention Analysis → Predictive Entity Verification”. As an example from US-Cities in Table 14, the triplet to be predicted is (Grand Saline, commute_time_comp, ?), for this question, there is only one city as answer named Fort Myers. When considering this comparative relation commute_time_comp, NumCoKE is able to focus more attention on the commute_time attribute associated with commute_time_comp, assigning 0.291 attention (the scores are obtained using Eq. (5) and the sum of all attributes is 1). However, RAKGE (DBLP:conf/kdd/KimKKPJP23) does not do so, paying more attention on other attributes, with only a fourth place ranking for commute_time with 0.0732 attention score. As a result, NumCoKE can predicte the true tail entity, while RAKGE only ranked Fort Myers in the third place. Another example (Cedarville, ranking_comp, ?) reflects the same findings. This shows RAKGE cannot capture complete relational context for modeling as it does not consider semantic interactions between entities and attributes.